# One Model, Any Conjunctive Query: Graph Neural Networks for Answering Queries over Incomplete Knowledge Graphs

### v2

@@ C.1 Base large query generation
We begin the dataset generation by sampling base formulas, to be later converted into instances for the QAC and QAR benchmarks. During sampling, we use four hyperparameters: $n_{\text{hub}},n_{\text{min}},p_{\text{const}}\text{ and }p_{\text{out}}$, whose different values contribute to creating different benchmark splits. The process is visualized in Figure˜6. A single base query is sampled as follows:
@@
For formulas sampled from FB15k-237, we choose $n_{\min}=15$, while for NELL instances, $n_{\min}=12$, due to the sparsity of the knowledge graph. We consider three different choices of the parameters $n_{\text{hub}},p_{\text{const}}\text{ and }p_{\text{out}}$, resulting in three distinct splits, namely “3-hub”, “4-hub” and “5-hub”, and sample 1000 formulas of each type. Using an SQL engine [Raasveldt and Mühleisen, 2019], we then solve these queries with respect to both observable and unobservable knowledge graphs, discarding those with no hard answers. The parameter values for each split are presented in Table˜8.
@@
C.2 Query answer classification datasets
@@
C.3 Query answer retrieval datasets
@@ C.4 Evaluation protocol
Appendix D Link predictors
As mentioned in Section˜5.1, we incorporate a link predictor into our architecture, to address the problem of deducing facts not presented in the observable knowledge graph. We consider three different model types from the existing CQA literature: transductive knowledge graph embedding method ComplEx [Trouillon et al., 2016] used in QTO and FIT, inductive (on nodes) method NBFNet [Zhu et al., 2021] employed by GNN-QE, and inductive (on nodes and relations) knowledge graph foundation model Ultra [Galkin et al., 2024a], lying at the heart of UltraQuery.
D.1 ComplEx
@@
For training, we follow the relation prediction methodology, presented in Chen et al. [2021], evaluating the loss as a sum over all known facts $r(a,b)\in E(G)$ of three cross-entropy losses, marginalizing the head, the relation and the tail:
@@
where $\mathcal{L}_{reg}$ is a nuclear 3-norm Lacroix et al. [2018] regularization term and the marginal probabilities are evaluated as:
@@
where $\tau$ is a factor controlling the temperature of the applied softmax function. During training, we set $\tau=1$. For each dataset, the model is trained using the AdaGrad Duchi et al. [2011] optimizer with a learning rate $0.1$ for 500 epochs, and the checkpoint maximizing validation accuracy is chosen for testing.
@@
To match the definition of a link predictor from Section˜3, the uncalibrated scores $\chi(r,a,b)$ assigned by the ComplEx model $\chi$ need to be converted into probabilities $\rho_{\mathbb{C}}(r,a,b)={\mathbb{P}}(r(a,b)\in E(\tilde{G})|\chi)$. We follow the ideas used in QTO [Bai et al., 2023] and FIT Yin et al. [2024], and set them as proportional to the marginal probabilities $p_{\chi,\tau}(b|a,r)$. By definition, $p_{\chi,\tau}(\cdot|a,r)$ defines a distribution over $V(G)$:
@@
D.2 NBFNet
As the second studied predictor, we consider Neural Bellman-Ford Network [Zhu et al., 2021], constituting the main processing unit in GNN-QE [Zhu et al., 2022]. For the AnyCQ experiments, we reuse the NBFNet checkpoints obtained from training GNN-QE over the considered datasets. We follow the configurations from the original repository – models are trained for 10 epochs, processing 48,000 instances per epoch for the FB15k-237 training, and 24,000 samples per epoch for NELL, with Adam [Kingma and Ba, 2015] optimizer with learning rate 0.005. We validate 0.25 to be the optimal threshold for binarizing GNN-QE predictions, and apply it for the small-query QAC experiments. When testing AnyCQ with the underlying NBFNet models, we first binarize the output of the NBFNet $\nu$:
@@ D.2 NBFNet
D.3 Ultra
Finally, to test AnyCQ’s ability of inductive link prediction over unseen relation, we consider Ultra [Galkin et al., 2024a], a prominent knowledge graph foundation models, as the third studied predictor for zero-shot inference. For the AnyCQ experiments, we directly apply the 3g checkpoints from the original Ultra repository, which are pre-trained on FB15k-237 [Toutanova et al., 2015], WN18RR [Dettmers et al., 2018], and CoDEx Medium [Safavi and Koutra, 2020] for 10 epochs with 80,000 steps per epoch, with AdamW [Loshchilov and Hutter, 2019] using learning rate of 0.0005. Similarly to the methodology applied to NBFNet, we binarize the output of the Ultra model $\upsilon$ when equipping it to AnyCQ. In this case, following validation, we choose $t=0.4$ as the best threshold for small query QAC experiments, and a higher $t=0.9$ for formulas in large QAC and QAR splits.
As an additional baseline, we compare another state-of-the-art CQA method over unseen relation UltraQuery [Galkin et al., 2024b], which also utilizes Ultra as its link predictor. We validate that 0.2 is the best answer classification threshold for UltraQuery checkpoints provided ni the original repository, trained only on the CQA benchmark based on FB15k-237. We highlight that the results of UltraQuery on NELL are hence zero-shot inference, since NELL is not in the pretraining dataset of the evaluated checkpoints.
D.4 Incorporating the observable knowledge graph
@@
D.5 Combination with AnyCQ
@@
As mentioned in Section˜5.4, the AnyCQ framework can be equipped with any link predictor capable of predicting relations over the studied knowledge graph. For this reason, to ensure that our choice matches the most accurate setup, we validate the performance of the predictors described in previous subsections, and test AnyCQ combined with ComplEx-based predictor with FIT scaling (Section˜D.1), NBFNet (Section˜D.2) and Ultra (Section˜D.3). Following validation, we choose NBFNet to be equipped for AnyCQ evaluations in all main experiments (Table˜1 and Table˜2).
As an additional ablation study, we generate the test results of the remaining combinations. The results on the QAC task are presented in Table˜11, while the scores on QAR benchmarks are shown in Table˜12. For the small-query QAC splits, we notice that NBFNet and Ultra strongly outperform ComplEx on positive formulas (“2p”, “3p”, “ip”, “pi”), while struggling more with queries with negations (“inp”, “pin”). The observed drop in Ultra results on the “3p” split is due to the model predicting too many links to be true, limiting the guidance from PE labels (almost each entity can be a head/tail of each relation). For large query classification, NBFNet produces much better F1-scores, exceeding 50% on almost all splits, while the remaining predictors consistently score under 40%. The evaluations on the QAR benchmarks further justify the choice of NBFNet as the equipped link predictor for AnyCQ – it consistently outperforms ComplEx and Ultra, achieving the best results on most FB15k-237-QAR, and all NELL-QAR splits.
@@
Interestingly, we again point out that the used Ultra model has not been trained on the NELL dataset. Regardless, it manages to match the performance of the ComplEx-based predictor on NELL-QAC and NELL-QAR benchmarks. Combining this observation with our ablation of transferability of AnyCQ models between datasets (Section˜6.4, Table˜4), we can assume that similar results would be achieved when running the evaluation with AnyCQ model trained on FB15k-237. Such framework would then answer queries over NELL in a true zero-shot, fully inductive setting. In future work, we look forward to exploring combinations of AnyCQ search engines trained over broad, multi-dataset data, with fully inductive link predictors (like Ultra), to achieve foundation models capable of answering arbitrary queries over arbitrary, even unseen, knowledge graphs.
@@
Appendix E Extended evaluation over QAR processing times
We analyze the processing times and retrieval performance of AnyCQ and the SQL engine on the QAR benchmarks, focusing on how both systems scale with query complexity.
To understand the impact of the timeout threshold on SQL performance, we evaluate it on all QAR splits using three different time limits: 30, 60, and 120 seconds. Results are reported in Table˜13. We observe that the SQL engine’s recall on easy queries improves marginally as the timeout increases, particularly for queries with multiple free variables. However, its performance consistently deteriorates with increasing arity, even under extended time limits. In contrast, AnyCQ maintains strong performance across all splits and consistently outperforms SQL on retrieving easy answers to high-arity queries. Additionally, it is worth noting that AnyCQ is also capable of retrieving hard answers, which SQL fails to find under any timeout setting.
Table˜14 presents the average processing times per query under the 60-second timeout. We find that AnyCQ exhibits significantly more stable and predictable runtimes compared to SQL. By design, AnyCQ’s processing time is independent of the number of free variables or other structural properties of the query, and remains efficient even on the larger NELL-QAR benchmark. In contrast, the SQL engine shows a steep increase in average runtime as query arity increases, with many queries exceeding the timeout. Notably, for queries with arity of at least 2, AnyCQ outperforms the SQL engine on average across all FB15k-237-QAR splits, while remaining competitive for single-variable queries. This reflects SQL’s sensitivity to query structure and its inefficiency in evaluating high-arity or cyclic formulas.
Across both tables, AnyCQ consistently demonstrates superior scalability and robustness. It handles structurally complex queries with multiple free variables more efficiently than SQL and is uniquely able to retrieve both easy and hard answers. Meanwhile, the SQL engine is unreliable in this setting - its performance is heavily dependent on query structure and its success rate is strongly limited by the time budget. These results validate our motivation for introducing QAR and support AnyCQ as a reliable neural alternative to classical engines for complex query answering.
Table 14: The comparison of processing times (in seconds) on the QAR task. We report the minimum, maximum and average processing time per split, together with the standard deviation (sd) and the number of unanswered samples due to the time restriction $n_{exc}$.
Dataset Split Arity Model $\min$ $\max$ avg sd $n_{exc}$
FB15k-237-QAR 3-hub $k=1$ AnyCQ 12.96 30.74 19.70 2.91 0
SQL 0.66 60.00 11.14 17.20 31
$k=2$ AnyCQ 13.01 30.56 19.57 2.99 0
SQL 0.32 60.00 30.87 24.84 143
$k=3$ AnyCQ 13.05 30.53 19.11 2.74 0
SQL 0.33 60.00 43.77 22.99 235
4-hub $k=1$ AnyCQ 13.37 32.72 20.44 2.98 0
SQL 0.30 60.00 13.90 19.34 40
$k=2$ AnyCQ 13.36 32.78 20.78 3.10 0
SQL 0.37 60.00 26.63 24.49 115
$k=3$ AnyCQ 14.24 32.99 20.42 2.89 0
SQL 0.36 60.00 32.77 24.88 154
5-hub $k=1$ AnyCQ 13.43 36.24 20.73 3.08 0
SQL 0.27 60.00 11.98 18.40 37
$k=2$ AnyCQ 13.95 36.25 21.74 3.05 0
SQL 0.57 60.00 23.47 24.40 106
$k=3$ AnyCQ 15.96 29.04 21.33 2.58 0
SQL 0.26 60.00 28.83 25.39 131
NELL-QAR 3-hub $k=1$ AnyCQ 21.87 59.57 35.35 6.27 0
SQL 0.16 60.00 6.59 13.77 15
$k=2$ AnyCQ 22.20 59.65 35.33 7.09 0
SQL 0.13 60.00 20.94 23.91 69
$k=3$ AnyCQ 18.81 59.69 35.29 7.61 0
SQL 0.14 60.00 21.77 25.13 75
4-hub $k=1$ AnyCQ 19.15 58.00 36.46 6.93 0
SQL 0.14 60.00 7.79 15.59 24
$k=2$ AnyCQ 19.14 57.58 35.70 7.54 0
SQL 0.13 60.00 17.47 22.54 51
$k=3$ AnyCQ 23.13 56.66 36.04 6.63 0
SQL 0.12 60.00 19.49 23.99 62
5-hub $k=1$ AnyCQ 23.40 60.00 37.27 6.58 3
SQL 0.17 60.00 6.71 15.01 22
$k=2$ AnyCQ 25.91 60.00 36.97 6.63 2
SQL 0.16 60.00 14.62 20.66 38
$k=3$ AnyCQ 24.25 60.00 36.84 5.91 1
SQL 0.17 60.00 17.44 22.18 45

### v3

@@ D.1 Base large query generation
We begin the dataset generation by sampling base formulas, to be later converted into instances for the QAC and QAR benchmarks. During sampling, we use four hyperparameters: $n_{\text{hub}},n_{\text{min}},p_{\text{const}}\text{ and }p_{\text{out}}$, whose different values contribute to creating different benchmark splits. The process is visualized in Figure˜4. A single base query is sampled as follows:
@@
For formulas sampled from FB15k-237, we choose $n_{\min}=15$, while for NELL instances, $n_{\min}=12$, due to the sparsity of the knowledge graph. We consider three different choices of the parameters $n_{\text{hub}},p_{\text{const}}\text{ and }p_{\text{out}}$, resulting in three distinct splits, namely “3-hub”, “4-hub” and “5-hub”, and sample 1000 formulas of each type. Using an SQL engine [duckdb], we then solve these queries with respect to both observable and unobservable knowledge graphs, discarding those with no hard answers. The parameter values for each split are presented in Table˜8.
@@
D.2 Query answer classification datasets
@@
D.3 Query answer retrieval datasets
@@ D.4 Evaluation protocol
Appendix E Link predictors
As mentioned in Section˜5.1, we incorporate a link predictor into our architecture, to address the problem of deducing facts not presented in the observable knowledge graph. We consider three different model types from the existing CQA literature: transductive knowledge graph embedding method ComplEx [trouillon2016complex] used in QTO and FIT, inductive (on nodes) method NBFNet [nbfnet] employed by GNN-QE, and inductive (on nodes and relations) knowledge graph foundation model Ultra [galkin2023ultra], lying at the heart of UltraQuery.
E.1 ComplEx
@@
For training, we follow the relation prediction methodology, presented in [relation_prediction], evaluating the loss as a sum over all known facts $r(a,b)\in E(G)$ of three cross-entropy losses, marginalizing the head, the relation and the tail:
@@
where $\mathcal{L}_{reg}$ is a nuclear 3-norm [nuclear_norm] regularization term and the marginal probabilities are evaluated as:
@@
where $\tau$ is a factor controlling the temperature of the applied softmax function. During training, we set $\tau=1$. For each dataset, the model is trained using the AdaGrad [adagrad] optimizer with a learning rate $0.1$ for 500 epochs, and the checkpoint maximizing validation accuracy is chosen for testing.
@@
To match the definition of a link predictor from Section˜3, the uncalibrated scores $\chi(r,a,b)$ assigned by the ComplEx model $\chi$ need to be converted into probabilities $\rho_{\mathbb{C}}(r,a,b)={\mathbb{P}}(r(a,b)\in E(\tilde{G})|\chi)$. We follow the ideas used in QTO [qto] and FIT [fit], and set them as proportional to the marginal probabilities $p_{\chi,\tau}(b|a,r)$. By definition, $p_{\chi,\tau}(\cdot|a,r)$ defines a distribution over $V(G)$:
@@
E.2 NBFNet
As the second studied predictor, we consider Neural Bellman-Ford Network [nbfnet], constituting the main processing unit in GNN-QE [gnn-qe]. For the AnyCQ experiments, we reuse the NBFNet checkpoints obtained from training GNN-QE over the considered datasets. We follow the configurations from the original repository – models are trained for 10 epochs, processing 48,000 instances per epoch for the FB15k-237 training, and 24,000 samples per epoch for NELL, with Adam [adam] optimizer with learning rate 0.005. We validate 0.25 to be the optimal threshold for binarizing GNN-QE predictions, and apply it for the small-query QAC experiments. When testing AnyCQ with the underlying NBFNet models, we first binarize the output of the NBFNet $\nu$:
@@ E.2 NBFNet
E.3 Ultra
Finally, to test AnyCQ’s ability of inductive link prediction over unseen relation, we consider Ultra [galkin2023ultra], a prominent knowledge graph foundation models, as the third studied predictor for zero-shot inference. For the AnyCQ experiments, we directly apply the 3g checkpoints from the original Ultra repository, which are pre-trained on FB15k-237 [fb15k237], WN18RR [Dettmers2018FB], and CoDEx Medium [safavi-koutra-2020-codex] for 10 epochs with 80,000 steps per epoch, with AdamW [loshchilov2019decoupledweightdecayregularization] using learning rate of 0.0005. Similarly to the methodology applied to NBFNet, we binarize the output of the Ultra model $\upsilon$ when equipping it to AnyCQ. In this case, following validation, we choose $t=0.4$ as the best threshold for small query QAC experiments, and a higher $t=0.9$ for formulas in large QAC and QAR splits.
As an additional baseline, we compare another state-of-the-art CQA method over unseen relation UltraQuery [galkin2024ultraquery], which also utilizes Ultra as its link predictor. We validate that 0.2 is the best answer classification threshold for UltraQuery checkpoints provided ni the original repository, trained only on the CQA benchmark based on FB15k-237. We highlight that the results of UltraQuery on NELL are hence zero-shot inference, since NELL is not in the pretraining dataset of the evaluated checkpoints.
E.4 Incorporating the observable knowledge graph
@@
E.5 Combination with AnyCQ
@@
As mentioned in Section˜5.4, the AnyCQ framework can be equipped with any link predictor capable of predicting relations over the studied knowledge graph. For this reason, to ensure that our choice matches the most accurate setup, we validate the performance of the predictors described in previous subsections, and test AnyCQ combined with ComplEx-based predictor with FIT scaling (Section˜E.1), NBFNet (Section˜E.2) and Ultra (Section˜E.3). Following validation, we choose NBFNet to be equipped for AnyCQ evaluations in all main experiments (Table˜1 and Table˜2).
As an additional ablation study, we generate the test results of the remaining combinations. The results on the QAC task are presented in Table˜11, while the scores on QAR benchmarks are shown in Table˜12.
For the small-query QAC splits, we observe that NBFNet and Ultra strongly outperform ComplEx on positive formulas (“2p”, “3p”, “ip”, “pi”), while struggling more with queries involving negations (“inp”, “pin”). The observed drop in Ultra’s results on the “3p” split is due to the model predicting too many links to be true, assigning high confidence scores to a large fraction of candidate triples. This behavior effectively expands the plausible search space, making it exponentially harder for the RL policy to converge to correct assignments. When too many entities are deemed likely, both local (LE) and global (PE) supervision signals become less informative, degrading the quality of guidance.
Overall, the performance of AnyCQ depends strongly on the characteristics of the underlying link predictor. As shown in Table 12, predictors trained with classification objectives (e.g., NBFNet, Ultra) generally outperform embedding-based models such as ComplEx. However, in practice, predictors with higher precision and more selective scoring distributions provide stronger search guidance for AnyCQ. These results suggest that, for complex query answering, prioritizing precision over recall in link predictor training yields more stable and effective search behavior.
For large-query classification, NBFNet demonstrates the most robust performance, achieving F1 scores exceeding 50% on nearly all splits, while the remaining predictors consistently stay below 40%. The evaluations on the QAR benchmarks confirm this trend: NBFNet consistently outperforms both ComplEx and Ultra, achieving the best results on most FB15k-237 and all NELL splits.
@@
Transferability. Interestingly, we again point out that the used Ultra model has not been trained on the NELL dataset. Regardless, it manages to match the performance of the ComplEx-based predictor on NELL-QAC and NELL-QAR benchmarks. Combining this observation with our ablation of transferability of AnyCQ models between datasets (Section˜6.3, Table˜3), we can assume that similar results would be achieved when running the evaluation with AnyCQ model trained on FB15k-237. Such framework would then answer queries over NELL in a true zero-shot, fully inductive setting. In future work, we look forward to exploring combinations of AnyCQ search engines trained over broad, multi-dataset data, with fully inductive link predictors (like Ultra), to achieve foundation models capable of answering arbitrary queries over arbitrary, even unseen, knowledge graphs.
Appendix F Extended evaluation over QAR
Table 13: Average F1-scores on easy and hard QAR samples, where $k$ is the number of free variables.
Dataset Model 3-hub 4-hub 5-hub
Easy instances
FB15k-237-QAR SQL 94.0 75.9 41.8 77.1 93.2 83.5 68.9 83.7 92.1 77.7 76.1 82.7
AnyCQ 90.8 91.4 83.4 89.2 88.3 83.1 83.8 85.2 85.0 80.2 68.7 78.8
NELL-QAR SQL 97.2 81.7 71.3 88.6 94.8 81.7 77.6 87.6 94.1 88.0 78.0 89.0
AnyCQ 96.3 94.4 92.7 95.1 94.5 94.0 88.6 93.1 94.4 91.6 89.5 92.5
Hard instances
FB15k-237-QAR SQL 0 0 0 0 0 0 0 0 0 0 0 0
AnyCQ 20.9 14.5 20.0 18.5 16.2 13.1 12.2 13.8 28.8 19.5 14.8 20.9
NELL-QAR SQL 0 0 0 0 0 0 0 0 0 0 0 0
AnyCQ 13.1 13.1 7.7 11.1 17.7 16.8 15.2 16.5 16.8 15.0 11.0 14.3
Easy and hard F1-scores. To better understand the performance differences between AnyCQ and SQL, we conduct a split evaluation over two categories of queries: (i) those with answers present in the observable graph (easy instances) and (ii) those requiring inference beyond the observable graph (hard instances). Results are reported in Table˜13.
On easy queries with a single free variable, SQL outperforms AnyCQ, though the margin is modest ($\leq 8\%$ relative). As the number of free variables increases, AnyCQ consistently surpasses SQL. While a classical solver could in principle answer these queries given unlimited time, our evaluation threshold of 60 seconds leads SQL to time out on a non-trivial subset. This indicates that AnyCQ can efficiently handle queries that impose substantial computational costs on symbolic engines.
On hard queries with no observable answers, SQL fails entirely, yielding $0\%$ F1 across all splits. In contrast, AnyCQ uses its integrated link predictor to infer missing edges while reasoning over complex query structures, successfully recovering answers across all settings. These findings validate our claims about the efficiency and effectiveness of AnyCQ: it not only enables prediction in incomplete knowledge graphs but also serves as a practical query answering engine in scenarios where classical solvers struggle.
Computational complexity of AnyCQ. Let $Q=\exists\vec{y}.\Phi(\vec{y})$ be a conjunctive Boolean query over a knowledge graph $G$. Denote by $|Q|$ the number of literals in $Q$, let $h$ be the maximum arity of a literal in $Q$, and let $\vec{c}=(c_{0},...,c_{s})$ be the mentioned constants. Then, the corresponding computational graph contains $(|V(G)|+1)\cdot|\vec{y}|+2\cdot|\vec{c}|+|Q|$ vertices and at most $|V(G)|\cdot(|\vec{y}|+h\cdot|Q|)+h\cdot|\vec{c}|\cdot|Q|$ edges. Since AnyCQ processes this graph in linear time, the complexity of a single search step is:
Importantly, this complexity does not depend on the structure of the query graph of $Q$ and scales only linearly with the sizes of the input formula and the KG $G$.
We validate this linearity, evaluating the average AnyCQ search step time for queries of different sizes from the ‘3-hub’ splits. The results, presented in Figure˜6, indicate that the empirical performance matches the theoretical analysis. In particular, Figure˜6 b) and c) show that the processing time, divided by the number of variables $|\vec{y}|$ or the number of literals $|Q|$ in the input query, respectively, does not grow as the size of the query increases. We even notice a slight decreasing trend, which we attribute to efficient GPU accelerations. The difference between step times on FB15k-237-QAR and NELL-QAR remains consistent with the relative sizes of the underlying knowledge graphs.
@@ Appendix F Extended evaluation over QAR
Experiments over impact of the timeout threshold
To further understand the impact of the timeout threshold on SQL performance, we evaluate it on all QAR splits using three different time limits: 30, 60, and 120 seconds. Results are reported in Table˜14. We observe that the SQL engine’s recall on easy queries improves marginally as the timeout increases, particularly for queries with multiple free variables. However, its performance consistently deteriorates with increasing arity, even under extended time limits. In contrast, AnyCQ maintains strong performance across all splits and consistently outperforms SQL on retrieving easy answers to high-arity queries.